# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ 8.0.1. RSH structure of $\mathcal{O}(\mathcal{E}_{Y})$
Let $B_{1}:=\Gamma(\mathscr{M}_{1})$ and let $\rho_{1}:\Gamma(\mathscr{M}_{2})\to\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})$ be the restriction map. Define
@@
by $\varphi_{1}(\varsigma)(x):=\pi_{2}(\xi)(x)$ for $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ with $\pi_{1}(\xi)=\varsigma$. First the existence of such an element $\xi$ follows from the surjectivity of $\pi_{1}$ by Lemma 7.7. In order to see that $\varphi_{1}$ is well defined, we have to show that if $\pi_{1}(\xi_{1})=\pi_{1}(\xi_{2})=\varsigma$, then $\pi_{2}(\xi_{1})(x)=\pi_{2}(\xi_{2})(x)$ for all $x\in\overline{Y_{2}}\setminus Y_{2}$. Since $\varsigma=(\varsigma_{1},\dots,\varsigma_{K}):=\pi(\xi_{1})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ satisfies the boundary decomposition property by Theorem 8.6, we have from (23) that
@@
Therefore, $\varphi_{1}$ is well defined, and obviously it is a -homomorphism. Let
@@
As in [16, Theorem 11.3.19], one can continue this process to form an RSH subalgebra $B_{k}$ of $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{k})$ for all $k=3,\dots,K$ simultaneously showing that at each step $k$, every element in $B_{k}$ satisfies the boundary decomposition property (only) up to $k$.
@@
For $k=1,2,\ldots,K$, there is a recursive subhomogeneous algebra $B_{k}\subset\oplus_{i=1}^{k}\Gamma(\mathscr{M}_{i})$ such that $\oplus_{i=1}^{k}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))=B_{k}$.
@@
We prove the theorem by induction on $k$. First note that by [32, Proposition 1.7] , $B_{1}:=\Gamma(\mathscr{M}_{1})$ is a recursive subhomogeneous algebra and $\pi_{1}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})$ is surjective by Lemma 7.7.
@@
is surjective onto a recursive subhomogeneous algebra $B_{k-1}$. To define $B_{k}$ we need a -homomorphism
$\varphi_{k-1}:B_{k-1}\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}}).$
For $(b_{1},b_{2},\ldots,b_{k-1})\in B_{k-1}$, since $\oplus_{i=1}^{k-1}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))=B_{k-1}$, there is $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ with $\oplus_{i=1}^{k-1}\pi_{i}(\xi)=(b_{1},b_{2},\ldots,b_{k-1})$. Define
@@
By Theorem 8.6, $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}$ is determined by $\pi_{i}(\xi)=b_{i}$, $1\leq i<k$, namely if $\oplus_{i=1}^{k-1}\pi_{i}(\xi)=(b_{1},\dots,b_{k-1})=\oplus_{i=1}^{k-1}\pi_{i}(\xi^{\prime})$ for $\xi$, $\xi^{\prime}\in\mathcal{O}(\mathcal{E}_{Y})$, then $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}=\pi_{k}(\xi^{\prime})|_{\overline{Y_{k}}\setminus Y_{k}}$. Hence $\varphi_{k-1}:B_{k-1}\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})$ is a well-defined -homomorphism. Define $B_{k}$ as
@@
where $\rho_{k-1}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})$ is the restriction. Then $\oplus_{i=1}^{k}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))$ is a $C^{*}$-subalgebra of the recursive subhomogeneous algebra $B_{k}$.
@@
By induction hypothesis, there is $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ such that
@@
Then the last coordinate $b_{k}-\pi_{k}(\xi)$ is not necessarily 00 over the whole $\overline{Y_{k}}$ but equal to 00 on $\overline{Y_{k}}\setminus Y_{k}$ since $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}=\varphi_{k-1}(b_{1},b_{2},\ldots,b_{k-1})=\rho_{k-1}(b_{k})$. It is thus enough to claim that for any $\varsigma\in\Gamma(\mathscr{M}_{k})$ with $\varsigma|_{\overline{Y_{k}}\setminus Y_{k}}=0$, there is $\eta\in\mathcal{O}(\mathcal{E}_{Y})$ with
@@
of $X$. We will define a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ such that $\eta:=\psi^{-1}(\widetilde{\xi})$ satisfies the claim, where $\psi$ is the isomorphism of Proposition 3.2. For this, we first define a local section $\xi:\mathcal{C}\to\mathscr{V}^{(m)}|_{\mathcal{C}}$ and then extend it to $\widetilde{\xi}$ over the whole space $X$. Since the closed subsets $\alpha^{i}(Y_{[l,k]})$, $11\leq l\leq j_{m}+$ 1, $r_{l-1}\leq i\leq r_{l}-1$, of $\mathcal{C}$ have overlaps (and contain $\alpha^{i}(\overline{Y_{k}})$), we need $\xi$ to be well-defined on any possible overlaps.
@@
for some $z_{1},z_{2}\in\overline{Y_{k}}$ and $i<j$. Then $\alpha^{j-i}(z_{2})=z_{1}\in Y$ while $0<j-i<r_{k}$. Thus $z_{2}\notin Y_{k}$, and so $z_{2}\in\overline{Y_{k}}\setminus Y_{k}$. Then $\varsigma^{(j)}\circ\alpha^{-j}(x)=\varsigma^{(j)}(z_{2})=0$. Also $\alpha^{r_{k}-j+i}(z_{1})=\alpha^{r_{k}}(z_{2})\in Y$ implies $z_{1}\notin Y_{k}$ since $0<r_{k}-j+i<r_{k}$. Then $z_{1}\in\overline{Y_{k}}\setminus Y_{k}$ and similarly we have $\varsigma^{(i)}\circ\alpha^{-i}(x)=\varsigma^{(i)}(z_{1})=0$. Thus $\xi$ is well-defined (to be 00) on overlaps of the form $\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{k}})$. To show that $\xi$ is well defined to be 00 on an overlap
@@
for $0\leq i\leq r_{k}-1-m$, suppose that $x=\alpha^{i}(\overline{y_{k}})\in\cup_{j=1}^{m}\alpha^{-j}(Y)$ for some $\overline{y_{k}}\in\overline{Y_{k}}$. Then $\alpha^{j+i}(\overline{y_{k}})\in Y$ for some $j$, but $10\leq j+i\leq r_{k}-$ 1 hence $\overline{y_{k}}\notin Y_{k}$. Then $\overline{y_{k}}\in\overline{Y_{k}}\setminus Y_{k}$ again, and then $\varsigma^{(i)}\circ\alpha^{-i}(x)=\varsigma^{(i)}(\overline{y_{k}})=0$. Now suppose $\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(Y_{k})\neq\emptyset$ for some $1\leq q<k$, $0\leq p<r_{q}$, $0\leq i<r_{k}$, with an element $x=\alpha^{p}(\overline{y_{q}})=\alpha^{i}(y_{k})$ where $\overline{y_{q}}\in\overline{Y_{q}}$ and $y_{k}\in Y_{k}$. Then $p\geq i$ by (26), and $\alpha^{r_{q}-p+i}(y_{k})=\alpha^{r_{q}-p}(x)=\alpha^{r_{q}-p}(\alpha^{p}(\overline{y_{q}}))=\alpha^{r_{q}}(\overline{y_{q}})\in Y$, which is not possible because $0<r_{q}-p+i<r_{k}$ and $y_{k}\in Y_{k}$. Thus $\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(\overline{Y_{k}})=\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(\overline{Y_{k}}\setminus Y_{k})$ on which $\varsigma^{(i)}\circ\alpha^{-i}=0$. Thus $\xi$ is well-defined over the closed subset $\mathcal{C}$ of $X$. Using the Tietze extension theorem for vector bundles, we further extend the domain of $\xi$ to the whole $X$ and get a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ over $X$ such that
@@
for $0\leq i\leq r_{k}-1-m$, and $\widetilde{\xi}|_{\cup_{j=1}^{m}\alpha^{-j}(Y)}=0$. Under the isomorphism $\psi:\mathcal{E}^{\otimes m}=\Gamma(\mathscr{V},\alpha)^{\otimes m}\to\Gamma(\mathscr{V}^{(m)},\alpha^{m})$, the property $\widetilde{\xi}|_{\cup_{j=1}^{m}\alpha^{-j}(Y)}=0$ implies that $\eta=\psi^{-1}(\widetilde{\xi})\in\mathcal{E}_{Y}^{\otimes m}=E_{m}\subset\mathcal{O}(\mathcal{E}_{Y})$. If $m=0$, then $\cup_{j=1}^{m}\alpha^{-j}(Y):=\emptyset$, $\mathcal{E}^{\otimes 0}=\Gamma(\mathscr{V}^{(0)},\alpha^{0})=C(X)$ and $\psi$ is the identity -homomorphism of $C(X)$.
@@
    Let $y\in X$ be a single point, and let $Y_{1}\supset Y_{2}\supset Y_{3}\supset\cdots$ be a sequence of closed subsets of $X$, each with non-empty interiors, satisfying
    $\cap_{n\in\mathbb{Z}_{>0}}Y_{n}=\{y\}.$
    Then
    $\mathcal{O}(\mathcal{E}_{\{y\}})\cong\underset{\longrightarrow}{\lim}\mathcal{O}(\mathcal{E}_{Y_{n}})$
    and so, by the above, $\mathcal{O}(\mathcal{E}_{\{y\}})$ is an approximately subhomogeneous (ASH) $\mathrm{C}^{*}$-algebra. The topological dimension of the RSH decomposition of $\mathcal{O}(\mathcal{E}_{Y_{n}})$ is bounded by the dimension of $X$. Thus if $X$ is finite dimensional, $\mathcal{O}(\mathcal{E}_{\{y\}})$ is an ASH with no dimension growth. It follows that $\mathcal{O}(\mathcal{E}_{Y_{n}})$ is $\mathcal{Z}$-stable and thus classifiable. Since $\mathcal{O}(\mathcal{E}_{Y_{n}})$ is a centrally large subalgebra of $\mathcal{O}(\mathcal{E})$, it too is $\mathcal{Z}$-stable and hence classifiable. Thus we recover Theorem 6.18 of [2].
@@
    If $(X,\alpha)$ is only assumed to have mean dimension zero (which means $\dim X$ may be infinite), the techniques above do not apply. However, using the RSH decomposition of $\mathrm{C}^{*}(C(X),C_{0}(X\setminus Y)u)\subset C(X)\rtimes_{\alpha}X$ for $Y\subset X$ a closed subset with non-empty interior, Elliott and Niu show that for any point $y\in X$ $\mathrm{C}^{*}(C(X),C_{0}(X\setminus\{y\})u)\subset C(X)\rtimes_{\alpha}X$ can be locally approximated by subhomogeneous $\mathrm{C}^{*}$-algebras with arbitrarily slow dimension ratio [15], which is enough to imply $\mathcal{Z}$-stability (and hence classifiability) of both $\mathrm{C}^{*}(C(X),C_{0}(X\setminus\{y\})u)$ and $C(X)\rtimes_{\alpha}X$.

### v2

@@ 8.0.1. An RSH structure for $\mathcal{O}(\mathcal{E}_{Y})$
Let $B_{1}:=\Gamma(\mathscr{M}_{1})$ and let $\rho_{1}:\Gamma(\mathscr{M}_{2})\to\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})$ be the restriction map. By Lemma 7.8, the map $\pi_{1}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})$ is surjective. Thus, for any $\varsigma_{1}\ \in\Gamma(\mathscr{M}_{1})$, there exists $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ such that $\pi_{1}(\xi)=\varsigma_{1}$. Define
@@
by
$\varphi_{1}(\varsigma_{1})(x):=\pi_{2}(\xi)(x).$
We claim that $\varphi_{1}$ is a well-defined -homomorphism. Suppose that $\pi_{1}(\xi_{1})=\pi_{1}(\xi_{2})=\varsigma_{1}$. By Theorem 8.6, $\varsigma=(\varsigma_{1},\dots,\varsigma_{K}):=\pi(\xi_{1})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ satisfies the boundary decomposition property. Then, using (24), we have that
@@
It follows $\varphi_{1}$ is well defined, and obviously it is a -homomorphism, which proves the claim. Let
@@
As in [18, Theorem 11.3.19], one can continue this process to form an RSH subalgebra $B_{k}$ of $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{k})$ for all $k=3,\dots,K$ simultaneously showing that at each step $k$, every element in $B_{k}$ satisfies the boundary decomposition property (only) up to $k$.
@@
For every $k=1,2,\ldots,K$, $B_{k}=\oplus_{i=1}^{k}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))$ is a recursive subhomogeneous $\mathrm{C}^{*}$-subalgebra of $\oplus_{i=1}^{k}\Gamma(\mathscr{M}_{i})$ such that $B_{1}=\Gamma(\mathscr{M}_{1})$, and for $k=2,\dots,K$,
$B_{k}=B_{k-1}\oplus_{\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})}\Gamma(\mathscr{M}_{k}).$
@@
We prove the theorem by induction on $k$. First note that by [35, Proposition 1.7] , $B_{1}:=\Gamma(\mathscr{M}_{1})$ is a recursive subhomogeneous algebra and $\pi_{1}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})$ is surjective by Lemma 7.8.
@@
is surjective onto the recursive subhomogeneous algebra $B_{k-1}$. We now construct $B_{k}$. For $(b_{1},b_{2},\ldots,b_{k-1})\in B_{k-1}$, since $\oplus_{i=1}^{k-1}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))=B_{k-1}$, there is $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ with $\oplus_{i=1}^{k-1}\pi_{i}(\xi)=(b_{1},b_{2},\ldots,b_{k-1})$. Define
@@
By Theorem 8.6, $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}$ is determined by $\pi_{i}(\xi)=b_{i}$, $1\leq i<k$, namely if $\oplus_{i=1}^{k-1}\pi_{i}(\xi)=(b_{1},\dots,b_{k-1})=\oplus_{i=1}^{k-1}\pi_{i}(\xi^{\prime})$ for $\xi$, $\xi^{\prime}\in\mathcal{O}(\mathcal{E}_{Y})$, then $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}=\pi_{k}(\xi^{\prime})|_{\overline{Y_{k}}\setminus Y_{k}}$. Hence
$\varphi_{k-1}:B_{k-1}\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})$
is a well-defined -homomorphism.
Define $B_{k}$ to be
@@
where $\rho_{k-1}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})$ is the restriction map. Then $\oplus_{i=1}^{k}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))$ is a $\mathrm{C}^{*}$-subalgebra of the recursive subhomogeneous algebra $B_{k}$.
@@
By the induction hypothesis, there is $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ such that
@@
The last coordinate $b_{k}-\pi_{k}(\xi)$ is not necessarily 00 over the whole $\overline{Y_{k}}$ but equal to 00 on $\overline{Y_{k}}\setminus Y_{k}$ since $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}=\varphi_{k-1}(b_{1},b_{2},\ldots,b_{k-1})=\rho_{k-1}(b_{k})$. It is thus enough to claim that for any $\varsigma\in\Gamma(\mathscr{M}_{k})$ with $\varsigma|_{\overline{Y_{k}}\setminus Y_{k}}=0$, there is $\eta\in\mathcal{O}(\mathcal{E}_{Y})$ with
@@
of $X$.
We will define a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ such that $\eta:=\psi^{-1}(\widetilde{\xi})$ satisfies the claim, where $\psi$ is the isomorphism of Proposition 3.2. For this, we first define a local section $\xi:\mathcal{C}\to\mathscr{V}^{(m)}|_{\mathcal{C}}$ and then extend it to $\widetilde{\xi}$ over the whole space $X$. Since the closed subsets $\alpha^{i}(Y_{[l,k]})$, $11\leq l\leq j_{m}+$ 1, $r_{l-1}\leq i\leq r_{l}-1$, of $\mathcal{C}$ have overlaps (and contain $\alpha^{i}(\overline{Y_{k}})$), we need $\xi$ to be well defined on any possible overlaps.
@@
for some $z_{1},z_{2}\in\overline{Y_{k}}$ and $i<j$. Then $\alpha^{j-i}(z_{2})=z_{1}\in Y$ while $0<j-i<r_{k}$. Thus $z_{2}\notin Y_{k}$, and so $z_{2}\in\overline{Y_{k}}\setminus Y_{k}$. Then $\varsigma^{(j)}\circ\alpha^{-j}(x)=\varsigma^{(j)}(z_{2})=0$. Also $\alpha^{r_{k}-j+i}(z_{1})=\alpha^{r_{k}}(z_{2})\in Y$ implies $z_{1}\notin Y_{k}$ since $0<r_{k}-j+i<r_{k}$. Then $z_{1}\in\overline{Y_{k}}\setminus Y_{k}$ and similarly we have $\varsigma^{(i)}\circ\alpha^{-i}(x)=\varsigma^{(i)}(z_{1})=0$. Thus $\xi$ is well defined (to be 00) on overlaps of the form $\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{k}})$. To show that $\xi$ is well defined to be 00 on an intersection of the form
@@
for $0\leq i\leq r_{k}-1-m$, suppose that $x=\alpha^{i}(\overline{y_{k}})\in\cup_{j=1}^{m}\alpha^{-j}(Y)$ for some $\overline{y_{k}}\in\overline{Y_{k}}$. Then $\alpha^{j+i}(\overline{y_{k}})\in Y$ for some $j$, but $10\leq j+i\leq r_{k}-$ 1 hence $\overline{y_{k}}\notin Y_{k}$. Then $\overline{y_{k}}\in\overline{Y_{k}}\setminus Y_{k}$ again, and then $\varsigma^{(i)}\circ\alpha^{-i}(x)=\varsigma^{(i)}(\overline{y_{k}})=0$. Now suppose $\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(Y_{k})\neq\emptyset$ for some $1\leq q<k$, $0\leq p<r_{q}$, $0\leq i<r_{k}$, with an element $x=\alpha^{p}(\overline{y_{q}})=\alpha^{i}(y_{k})$ where $\overline{y_{q}}\in\overline{Y_{q}}$ and $y_{k}\in Y_{k}$. Then $p\geq i$ by (27), and $\alpha^{r_{q}-p+i}(y_{k})=\alpha^{r_{q}-p}(x)=\alpha^{r_{q}-p}(\alpha^{p}(\overline{y_{q}}))=\alpha^{r_{q}}(\overline{y_{q}})\in Y$, which is not possible because $0<r_{q}-p+i<r_{k}$ and $y_{k}\in Y_{k}$. Thus $\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(\overline{Y_{k}})=\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(\overline{Y_{k}}\setminus Y_{k})$ on which $\varsigma^{(i)}\circ\alpha^{-i}=0$. Thus $\xi$ is well defined over the closed subset $\mathcal{C}$ of $X$.
Using the Tietze extension theorem for vector bundles, we further extend the domain of $\xi$ to the whole $X$ to obtain a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ over $X$ such that
@@
for $0\leq i\leq r_{k}-1-m$, and such that
$\widetilde{\xi}|_{\cup_{j=1}^{m}\alpha^{-j}(Y)}=0.$
Under the isomorphism $\psi:\mathcal{E}^{\otimes m}=\Gamma(\mathscr{V},\alpha)^{\otimes m}\to\Gamma(\mathscr{V}^{(m)},\alpha^{m})$, the property $\widetilde{\xi}|_{\cup_{j=1}^{m}\alpha^{-j}(Y)}=0$ implies that $\eta=\psi^{-1}(\widetilde{\xi})\in\mathcal{E}_{Y}^{\otimes m}\cong E_{m}\subset\mathcal{O}(\mathcal{E}_{Y})$. If $m=0$, then $\cup_{j=1}^{m}\alpha^{-j}(Y):=\emptyset$, $\mathcal{E}^{\otimes 0}=\Gamma(\mathscr{V}^{(0)},\alpha^{0})=C(X)$ and $\psi$ is the identity -homomorphism of $C(X)$.
@@
Observe that the structure of the matrix bundles $\mathscr{M}_{k}$ in the iterated pullback
$\Gamma(\mathscr{M}_{1})\oplus_{\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})}\oplus\dots\oplus_{\Gamma(\mathscr{M}_{K}|_{\overline{Y_{K}}\setminus Y_{K}})}\Gamma(\mathscr{M}_{K}),$ (28)
is dictated by the line bundle $\mathscr{V}$. In particular, if $\mathscr{V}$ is trivial, then so too are the $\mathscr{M}_{k}$. In that case $\Gamma(\mathscr{M}_{k})\cong C(\overline{Y_{k}},M_{r_{k}})$ and we obtain the same decomposition as the one given in (8).
In the case of a non-trivial line bundle, we arrive at the following.
Theorem 8.8.
Let $X$ be an infinite compact metric space, $\alpha:X\to X$ a minimal homeomorphism, $\mathscr{V}$ a line bundle over $X$ and $Y\subset X$ a closed subset with non-empty interior. Let $r_{1}<\dots<r_{K}$ denote the distinct first return times to $Y$. The orbit breaking algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ has a recursive subhomogeneous decomposition $B$ such that
  1. (1)
    the length $l$ of $B$ is at least $K$,
  2. (2)
    the base spaces $Z_{1},\dots,Z_{l}$ have dimension bounded by $X$, and there is $l_{0}=0<l_{1}<l_{2}<\dotsm<l_{K}=l$ such that
    $\bigcup_{j=l_{k-1}+1}^{l_{k}}Z_{j}=\overline{\{y\in Y\mid r(y)=r_{k}\}}$
    for every $1\leq k\leq K$,
  3. (3)
    the matrix sizes are exactly $r_{1},\dots,r_{K}$.
Proof.
Let $\mathcal{E}_{Y}:=C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha)$. By the previous theorem, we have that
By [35, Proposition 1.7], $\Gamma(\mathscr{M}_{k})$, for every $1\leq k\leq K$, is an RSH algebra with matrix sizes $r_{k}$ with base spaces $Z_{l_{k-1}+1},\dots,Z_{l_{k}}$ for some $l_{k}\geq l_{k-1}+1$ whose union is the base space of the bundle, that is
$\bigcup_{j=l_{k-1}+1}^{l_{k}}Z_{j}=\overline{\{y\in Y\mid r(y)=r_{k}\}},$
It follows from [35, Proposition 3.2] that $\mathcal{O}(\mathcal{E}_{Y})$ itself has an RSH decomposition with length at least $K$, whose bases spaces are bounded in dimension by the dimension of $X$ and satisfy (2), and whose matrix sizes are $r_{1},\dots,r_{K}$. ∎
A main motivation for the RSH construction is to provide classification results for the $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles and their simple orbit-breaking subalgebras. In [17], the authors will show that classification is possible under the assumption that the homeomorphism has mean dimension zero. However, Theorem 8.8 is already enough to give an alternative proof of the classification results from [2], which is to say, the case where $\dim X<\infty$.
Let $A$ be a simple, separable, nuclear, unital $\mathrm{C}^{*}$-algebra. The Elliott invariant of $A$, denoted $\textup{Ell}(A)$ is given by
$\textup{Ell}(A)=(K_{0}(A),K_{0}(A)_{+},[1_{A}],K_{1}(A),T(A),\rho),$
where $(K_{0}(A),K_{0}(A)_{+},[1_{A}],K_{1}(A))$ is the (pointed, ordered) $K$-theory of $A$, $T(A)$ the tracial state simplex, and $\rho:K_{0}(A)\times T(A)\to\mathbb{R}$ the pairing map defined by $\rho([p]-[q],\tau)=\tau(p)-\tau(q)$, for $\tau$ the (non-normalized) inflation of a tracial state to a suitable matrix algebra over $A$.
Corollary 8.9 (cf. [2, Theorem 6.18]).
Let $X$ and $Y$ be finite-dimensional infinite compact metric spaces. Let $\mathcal{E}=\Gamma(\mathscr{V},\alpha)$ and $\mathcal{F}=\Gamma(\mathscr{W},\beta)$ be a Hilbert $C(X)$-module and $C(Y)$-module, respectively, with respect to the line bundles $\mathscr{V}$ over $X$, $\mathscr{W}$ over $Y$, and minimal homeomorphisms $\alpha:X\to X$, $\beta:Y\to Y$. Then
$\mathcal{O}_{C(X)}(\mathcal{E})\cong\mathcal{O}_{C(Y)}(\mathcal{F})$
if and only if
$\textup{Ell}(\mathcal{O}_{C(X)}(\mathcal{E}))\cong\textup{Ell}(\mathcal{O}_{C(Y)}(\mathcal{F})).$
Proof.
Let $y\in X$ be a single point, and let $Y_{1}\supset Y_{2}\supset Y_{3}\supset\cdots$ be a sequence of closed subsets of $X$, each with non-empty interior, satisfying
$\cap_{n\in\mathbb{Z}_{>0}}Y_{n}=\{y\}.$
Then
$\mathcal{O}_{C(X)}(\mathcal{E}_{\{y\}})\cong\underset{\longrightarrow}{\lim}\,\mathcal{O}_{C(X)}(\mathcal{E}_{Y_{n}}).$
By Theorem 8.8, the orbit-breaking algebras $\mathcal{O}(\mathcal{E}_{Y_{n}})$ are recursive subhomogeneous $\mathrm{C}^{*}$-algebras whose base spaces are bounded in dimension by the dimension of $X$. By [45, Theorem 6.1], it follows that $\mathcal{O}_{C(X)}(\mathcal{E}_{Y_{n}})$ has decomposition rank bounded by $\dim X$. Since $\mathcal{O}_{C(X)}(\mathcal{E}_{\{y\}})$ is the inductive limit of the $\mathcal{O}_{C(X)}(\mathcal{E}_{Y_{n}})$, the decomposition rank $\mathcal{O}_{C(X)}(\mathcal{E}_{\{y\}})$ is also bounded by $\dim X$ [27, 3.3 (ii)]. It follows that $\mathcal{O}_{C(X)}(\mathcal{E}_{\{y\}})$ is $\mathcal{Z}$-stable [46, Theorem 5.1]. Since $\mathcal{O}(\mathcal{E}_{\{y\}})$ is a centrally large subalgebra of $\mathcal{O}_{C(X)}(\mathcal{E})$ [2, Theorem 6.16], it too is $\mathcal{Z}$-stable. Similarly, $\mathcal{O}_{C(Y)}(\mathcal{F})$ is $\mathcal{Z}$-stable. Thus $\mathcal{O}_{C(X)}(\mathcal{E})$ and $\mathcal{O}_{C(Y)}(\mathcal{F})$ are simple, separable, unital, nuclear, $\mathcal{Z}$-stable $\mathrm{C}^{*}$-algebras and the result follows from [15]. ∎
@@
    If the bundle $\mathscr{V}$ is trivial, then the RSH decomposition of Theorem 8.8 is the same as the RSH decomposition for $\mathrm{C}^{*}(X,C_{0}(X\setminus Y)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$ given in [30].
@@
    If $(X,\alpha)$ is only assumed to have mean dimension zero (which means $\dim X$ may be infinite), we can’t deduce finite decomposition rank of $\mathcal{O}(\mathcal{E}_{\{y\}})$, which was used in the proof of Corollary 8.9 to show classification. This is because the base spaces of the RSH decomposition for each $\mathcal{O}(\mathcal{E}_{Y_{n}})$ can be infinite dimensional. However, using the RSH decomposition of $C^{*}(C(X),C_{0}(X\setminus Y)u)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$ for $Y\subset X$ a closed subset with non-empty interior, Elliott and Niu show that for any point $y\in X$, the orbit-breaking algebra $C^{*}(C(X),C_{0}(X\setminus\{y\})u)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$ can be locally approximated by subhomogeneous $\mathrm{C}^{*}$-algebras with arbitrarily small dimension ratio [16], which is enough to imply $\mathcal{Z}$-stability (and hence classifiability) of both the orbit-breaking algebra $C^{*}(C(X),C_{0}(X\setminus\{y\})u)$ and its containing crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$. In [17], we will use the RSH construction of Theorem 8.8 to generalize this to the Cuntz–Pimsner algebras $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(\Gamma(\mathscr{V},\alpha)_{\{y\}})$.